OpenStudy (anonymous):

The sum of 4 consecutive integers is 134, what is the product of the first and second?

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

guess and check do you know what "consecutive" means?

OpenStudy (anonymous):

I used to, like its been so long since ive used it :P

OpenStudy (anonymous):

i means one and then the next one, like 20 and 21 or 32 and 33

OpenStudy (anonymous):

we can use an equation to solve, but it is probably easier just to figure it out by guessing i guess \(30+31+32+33\) but that is not right, because \(30+31+32+33=126\) so i started too low

OpenStudy (anonymous):

SO would i write X+X+1+x+2+X+3=134 and then when i answer it i add the first and second?

OpenStudy (anonymous):

yes after you write that you might write \[4x+6=134\]

OpenStudy (anonymous):

subtract 6 divide by 4 to get the first number \(x\)

OpenStudy (anonymous):

Ohh ok! thats 32!

OpenStudy (anonymous):

like x would be 32

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then do I add the 1 and 2 and 3 to X and add those up?

OpenStudy (anonymous):

actually i think you are asked for the product of the first two since the first one is 32 you know what the second one is right?

OpenStudy (anonymous):

Its 33 right? and then do you add the 32 and 33?

OpenStudy (anonymous):

no, "product" means \(\times\)

OpenStudy (anonymous):

like "the product of 6 and 5 is 30"

OpenStudy (anonymous):

but yes, the second one is 33, so your job now is to compute \(32\times 33\)

OpenStudy (anonymous):

Ohhhhhhhhhhh! ok! and thats 65! Thanks!!!!

OpenStudy (anonymous):

no no times, not plus!!

OpenStudy (anonymous):

\(32\times 33\) not \(32+33\)

OpenStudy (anonymous):

Wait what? i thought you added?

OpenStudy (anonymous):

no "product" means "times"

OpenStudy (anonymous):

the product of \(a\) and \(b\) is \(a\times b\)

OpenStudy (anonymous):

question says " what is the product of the first and second?"

OpenStudy (anonymous):

not the sum of the first and second

OpenStudy (anonymous):

psh oh haha ok, well my calculator says its 1056?

OpenStudy (anonymous):

so does mine

OpenStudy (anonymous):

now i have a question i know what a swimmer is, but what is a gaswimmer?

OpenStudy (anonymous):

Haha well i wanted the name swimmer4life but it was taken, so Gaswimmer is basically (Georgia)Swimmer because im from georgia :P haha

OpenStudy (anonymous):

oooh i see

OpenStudy (anonymous):

Haha yeah its been confusing people, are you in High school?

OpenStudy (anonymous):

oh no, not for a long time

OpenStudy (anonymous):

haha i was wondering because you explained it really well, alot better than my teacher -_- i just started algebra 1 and like its giving me the biggest pain :P but thanks a ton for the help!!!

OpenStudy (anonymous):

yw if you have any more questions, just post!

OpenStudy (anonymous):

Is this correct? (You dont have to do it I was just wondering?) 72m/h = Ft per second